# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

### v1

@@ 6.2 Possible relevance for observations and experiments
In this article, we focus only on revealing the novel scenario and mechanism of generating UHE particles, and pay no attention to quantitatively discussing this scenario compared with the observations of UHE and VHE particles. Nevertheless, we mention some observationally and experimentally relevant features below:
@@
    Interacting with magnetic fields, soft photons and other low-energy particles around sources, these UHE electrons and protons from the opaque fluid spot of highly energetic photons and pairs produce VHE photons and neutrinos to TeV or PeV energy scales. In addition to astrophysical sources, such an energetic spot would be created in future by strong laser-beam experiments. The probability of producing UHE particles could be non-trivial for detection, if the strong laser field spot is not only sufficiently energetic $T_{\gamma}\gg m_{e}$ and dense $A\gg 1$, but also has a lifetime longer than the time scale $\tau^{e}_{{}_{\rm KN}}=(\sigma^{e}_{{}_{\rm KN}}n_{\gamma}c)^{-1}$ (4.2) for the runaway process.
@@
    We come to the GRBs’ cases. As shown in Fig. 7, the trapped fireballs (horizon halos) cool via continuous emission of UHE electrons and protons. Interacting with the surrounding fields and materials, the former mainly leads to VHE photons, and the latter to VHE neutrinos. Before outward expanding fireballs become optically thin, we probably cannot see VHE photons because of great opacity due to two-photon annihilation $\gamma\gamma\rightarrow e^{+}e^{-}$ of the Breit-Wheeler process (2.6). It strongly implies that the VHE photons most probably observed are after expanding fireballs become transparent, namely, VHE photon luminosity increases soon after the main GRBs’ bursts [24]. During the afterglow era, UHE protons $\gamma_{p}\gg 1$ emitted from the horizon halo continuously collide with the ejector of the Lorentz factor $\Gamma<\gamma_{p}$, which then emits VHE photons and neutrinos. Therefore, their light curves follow the law of the ejector’s kinematic motion in time, similarly to the external shock emission of the ejector colliding with the interstellar medium. However, their energy spectra become softer in time because the horizon halos cool down. The time and luminosity of UHE and VHE continuous emissions depend on the horizon-halo temperature $T_{\gamma}$ and acceleration number $A$, as well as the target (ejector) density.
@@
    The scenario and mechanism in principle apply to the early Universe at large particle density and high temperature $T_{\gamma}>2m_{e}$ before the Universe becomes transparent. Local dense and inhomogeneous radiation fields accelerate charged particles to high energies. It implies that in the cosmic background observed today, there should be imprints of excess UHE and VHE particles originating from the early Universe.
Further investigations in theoretical studies, numerical simulations, observations and experiments are necessary in future.

### v2

@@ 7.2 Possible relevance for observations and experiments
In this article, we focus only on revealing the novel scenario and mechanism of generating UHE particles, and pay no attention to quantitatively discussing this scenario compared with the observations of UHE and VHE particles. Nevertheless, we mention some observationally and experimentally relevant features, in particular for UHE and VHE particles accompanying GRB phenomena observed.
Interacting with magnetic fields, soft photons and other low-energy particles around sources, these UHE electrons and protons from the opaque radiation fluid spot of highly energetic photons and pairs produce VHE photons and neutrinos to TeV or PeV energy scales. We come to GRBs’ cases. As shown in Fig. 7, the trapped fireballs (horizon halos) cool via continuous emission of UHE electrons and protons. Interacting with the surrounding fields and materials, the former mainly leads to VHE photons, and the latter to VHE neutrinos. This scenario strongly implies the logical and causal reasoning features that are relevant to the observations of VHE photons accompanying GRBs’ phenomena:
@@
    UHE electrons (protons) and the subsequent VHE photons and neutrinos are continuous emissions in the fireshell lifetime. They should correlate with each other with the GRBs’ sources in observations.
@@
    Before outward expanding fireballs become optically thin, we probably cannot see VHE photons because of great opacity due to two-photon annihilation $\gamma\gamma\rightarrow e^{+}e^{-}$ of the Breit-Wheeler process (2.7). During the prompt emissions for bursts, the VHE photons most probably observed are after expanding fireballs become transparent, namely, VHE photon luminosity increases soon after the main GRBs’ bursts occur  [^8] The features (1) and (2) were pointed out in Ref. [17].
@@
    During the afterglow era, UHE protons $\gamma_{p}\gg 1$ emitted from the horizon fireshell continuously collide with the ejector of the Lorentz factor $\Gamma<\gamma_{p}$, which then emits VHE photons and neutrinos. Therefore, their (VHE) light curves follow the law of the ejector’s kinematic motion in time, similarly to the external shock emission of the ejector colliding with the interstellar medium. However, their (VHE) energy spectra become softer in time because the horizon fireshell cool down. The time and luminosity of UHE and VHE continuous emissions depend on the horizon fireshell temperature $T_{\gamma}$ and acceleration number $A$ (7.1), as well as the target (ejector) density.
In contrast, according to the usual modelling interpretation of the TeV emission in GRBs, ultra-relativistic electrons are accelerated by internal shocks of fireballs during the prompt emission or by external shocks between the ejector and interstellar medium (ISM) during the afterglow phase.
@@
In addition to astrophysical sources, the scenario and mechanism in principle apply to the early Universe at large particle density and high temperature $T_{\gamma}>2m_{e}$ before the Universe becomes transparent. Local dense and inhomogeneous radiation fields accelerate charged particles to high energies. It implies that in the cosmic background observed today, there should be imprints of excess UHE and VHE particles originating from the early Universe. Moreover, such an energetic spot would be created in future by strong laser-beam experiments in ground laboratories. The probability of producing UHE particles could be non-trivial for detection, if the strong laser field spot is not only sufficiently energetic $T_{\gamma}>10m_{e}$ and dense $A\gg 1$, but also has a lifetime longer than the time scale $\tau^{e}_{{}_{\rm KN}}=(\sigma^{e}_{{}_{\rm KN}}n_{\gamma}c)^{-1}$ (4.2) for the runaway process.
8 Summary and remarks
In this article, we study the possibility of how electrons and protons are accelerated to UHE particles in an opaque fire spot of dense photons and $e^{-}e^{+}$ pairs at a large temperature $T_{\gamma}\gg m_{e}$. We have revealed a possible acceleration mechanism for UHE charged particles, which mainly attribute to (i) the Compton-rocket effect of strong radiation force acting on electrons for acceleration; (ii) the instability and runaway dynamics occur when accelerated electrons undergo the KN scattering process, whose cross section decreases with increasing electron energy; (iii) such drifting electrons develop a local electric field acting on protons for acceleration. For qualitative analysis and theoretical illustrations, we use the Compton-rocket equation for electrons’ energy gain and the rate equation of electrons’ energy loss in Thomson and KN scatterings in a simplified one-dimensional model. As a result, we obtain the non-trivial probability of accelerating electrons and protons reaching a large Lorentz factor $\gamma_{e}\sim 10^{9}$ over a small distance $d\sim 10^{-2}$ cm.
As an astrophysical source for the opaque fire spots of dense photons and pairs at large temperatures, we consider the opaque fireballs formed in gravitational collapses. The outer parts of fireballs expand, accounting for the phenomena of GRBs. Whereas the inner parts of fireballs, fireshell, slowly contract and get dense near the horizon due to strong gravity and large pressure. The inner part of fireballs, fireshell as an energy source, possesses large opacity and thermal energy with a small amount $B$ of electrons and protons from baryon matter contamination, which are comparable with counterparts of the outer part of fireballs. Approximating the fireshell as a metastable state, we calculate the luminosities and spectra of UHE emission and blackbody radiation. In this scenario, we point out distinct features of produced UHE and VHE particles, which are relevant for observations and in contrast with the usual interpretations of the TeV emission in GRBs.
However, we emphasise that the systems and dynamics under investigation are rather complex, consisting of photons and pairs, protons and electrons, various interacting processes, whose times and lengths run from the microscopic to the macroscopic scales. In these preliminary studies, the models and conditions used to present the idea and scenario are simplified, the fraction $N_{e}/\bar{N}_{e}$ and the analyses used to show the results are qualitative. The one-dimensional electric field between drifting protons and electrons should be generalised to the three-dimensional case. Full interaction rates and cross sections depending on energy and momentum should be considered to investigate how the instability of electron perturbations develops in the strong radiation force of dense and energetic photons and pairs, leading to an avalanche runaway process. Needless to say, it is necessary to develop a numerical code for performing full simulations to verify the present scenario of accelerating electrons and protons to ultra-high energies. The dynamics are attributed to the radiation force in an opaque fire spot of dense photons and pairs at temperatures that exceed the electron mass and have a large spatial gradient in a preferential direction. Moreover, it is also necessary to study how such firespots can be established in high-energy astrophysical processes in strong gravitational environments, and determine their temperatures, sizes and lifetimes. In observations, it is worthwhile to phenomenologically examine the three distinct features of VHE photon emissions in GRBs mentioned in Sec. 7.2, and possible relations between observed cosmic ray events of UHE particles and VHE emissions of photons and neutrinos, and correlations with high-energy astrophysical sources.